OpenStudy (matimaticas):

what is the 27th root of x^9

OpenStudy (anonymous):

\[(x ^{9})^{1/27} = x ^{1/3} = \sqrt[3]{x}\]

OpenStudy (anonymous):

when you take nth root of an expression, it means 1/n th power of it

OpenStudy (matimaticas):

thx